OpenStudy (diana.xl):

Two models of cellular telephones, red and blue, are stored in boxes. One box weighs twelve pounds and contains four of the red models and one blue model. Another box weighs eight pounds and contains one blue model and two of the red models. How much does each model weigh? The blue model weighs two pounds, and the red model weighs four pounds. The red model weighs two pounds, and the blue model weighs six pounds. The red model weighs two pounds, and the blue model weighs four pounds. The blue model weighs five pounds, and the red model weighs eight pounds.

OpenStudy (diana.xl):

@imqwerty

imqwerty (imqwerty):

let the weight of red model be r let the weight of blue model be b according to the given information we have these equations-> 4r+b=12 and 4b+2r=8 from equation 1 4r+b=12 b=12-4r put b=12-4r in 2nd eq b+2r=8 12-4r+2r=8 12-8=2r 4=2r r=2pounds b=12-4r b=12-4(2) b=12-8 b=4pounds

OpenStudy (diana.xl):

my answer is c yea?

imqwerty (imqwerty):

yea B)

imqwerty (imqwerty):

the answer is C

OpenStudy (diana.xl):

Ty

imqwerty (imqwerty):

np B)